# Weighted Graph Structure Learning with Attention Denoising for Node Classification

### v1

@@ I Introduction
To enhance the ability of the model to handle noisy edges, we combine two approaches to address node classification in weighted graphs, facing two primary challenges. First, edge weights in weighted graphs can describe fine-grained relationships between nodes, with higher weights indicating stronger connections and greater influence on information propagation[20][21]. However, existing weight learning methods primarily focus on the structural characteristics of the graph and do not adequately account for edge weights. How to learn more fine-grained weights between nodes is a challenge. Second, noisy edges in weighted graphs are characterized by anomalous edge weights. Specific scenarios, such as network failures affecting communication between users for a period, can cause edge weights that describe communication frequency to deviate significantly from normal conditions[22]. How to adaptively correct anomalous edge weights and remove noisy edges is a challenge.
To address these challenges, we propose an Edge Weight-aware Graph Structure Learning (EWGSL) method, a novel model that employs edge weight-aware methods to learn fine-grained weights and employing Graph Structure Learning (GSL)[22] to reduce noise interference. Leveraging the excellent scalability of Graph Attention Networks (GATs)[23], we use GAT to learn node representations. We redefine the calculation of attention coefficients in GAT by incorporating both node features and edge weights to capture the relationships between nodes accurately. We apply GSL to sparsify the attention coefficients, setting the attention weights of irrelevant nodes to zero and concentrating them on relevant nodes to effectively identify noisy edges. Additionally, we modify the noise contrastive estimation (InfoNCE)[24] loss function to dynamically utilize the weights from the denoised graph structure, thereby enhancing node classification performance.
@@ II-A Weight Learning Methods
WFSM-MaxPWS[20] prunes the subgraph expansion process by calculating the ”Maximum Possible Weighted Support”, retaining only subgraphs with higher weighted support. This reduces computational overhead and improves mining efficiency while ensuring the completeness of the results. MetaGC[8] adaptively adjusts the weight of each edge in the graph using a meta-model, specifically lowering the weight of noisy edges to mitigate their impact on the clustering results. Additionally, it employs a decomposable clustering loss function and end-to-end training, allowing the method to enhance robustness against noisy edges without relying on external denoising steps.
@@ II-B Structure Learning Methods
Structure learning methods handle noisy edges by optimizing the graph structure, primarily through identifying and removing noisy edges, and reconstructing the graph’s connectivity. This enhances the propagation of information between similar nodes. This approach ensures a more accurate graph structure, improves the quality of node representations, and reduces the interference of noise on model learning.
@@ III-A Problem Definition
Problem Formulation: Given a weighted graph $\mathrm{G}$, and a small set of nodes $\mathrm{V_{L}}$ with labels $\mathrm{Y_{L}}$, the goal of node classification task is to predict unlabelled nodes $\mathrm{V_{U}}$ with $\mathrm{Y_{U}}$, where ${\mathrm{V_{L}}}\cup{\mathrm{V_{U}}}=\mathrm{V}$.
@@ III-B Edge-aware Weight Learning
$e_{ij}=\rho_{ij}\cdot{\boldsymbol{a}(\textbf{W}{\vec{h}_{i}},\textbf{W}{\vec{h}_{j}})},$ (1)
@@
$\rho_{ij}=\frac{w_{ij}}{\sum\limits_{k\in{{\mathcal{N}_{i}}}}w_{ik}}\;(i\notin{\mathcal{N}_{i}}),$ (2)
@@ III-C Graph Sparsity Structure Learning
${e}_{ij}^{\prime}=\mathrm{{\alpha}\text{-}entmax}({e_{ij}})={[(\alpha-1){e_{ij}}-\tau]_{+}^{1/{\alpha-1}}},$ (4)
@@
${\vec{h}_{i}}^{\prime}=\mathrm{F}(\Delta{\vec{h}_{i}}^{\prime})=\mathrm{F}(\beta_{k}{\vec{h}_{ki}}^{\prime})\quad(k=1,2,\ldots,K),$ (7)
@@
$\mathcal{H}=[{\vec{h}_{1}}^{\prime},{\vec{h}_{2}}^{\prime},\cdots,{\vec{h}_{n}}^{\prime}],$ (8)
@@ III-D Loss Function
$\mathcal{M}=[\mathrm{softmax}({\vec{h}_{1}}^{\prime}),\mathrm{softmax}({\vec{h}_{2}}^{\prime}),\cdots,\mathrm{softmax}({\vec{h}_{n}})].$ (9)
@@
$\hat{y}_{i}={\mathrm{{\underset{{\textit{k}}\in\{1,2,\cdots,c\}}{argmax}}}m_{ik}}.$ (10)
@@ 13      Backpropagate and update model parameters using $L$;
$\vspace{-0.2cm}L_{I}=-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\lambda_{p}\exp\left({\text{sim}({\vec{h}_{i}}^{\prime},{\vec{h}_{j}}^{\prime})}/{t}\right)}{{\sum\limits_{k\in N_{i}}}\lambda_{n}\exp\left({\text{sim}({\vec{h}_{i}}^{\prime},{\vec{h}_{k}}^{\prime})}/{t}\right)},\vspace{-0.2cm}$ (12)
@@ IV-A Experiment Settings
ML-100k__\__ES [^1] https://anonymous.4open.science/r/EWGSL-C021/ML-100k__\__ES/ is constructed from the open dataset MovieLens 100K [^2] https://grouplens.org/datasets/movielens/. The dataset includes multiple files, and we focus on u.data and u.item. We will release this dataset.
@@
We use a three-layer GAT to learn node representations, with hidden layers of sizes 256 and 128, and input and output sizes of $n$ and $c$, respectively [^3] https://anonymous.4open.science/r/EWGSL-C021/EWGSL/. The model trains for 100 epochs with a learning rate of 0.005. Based on the dataset construction, we divide the datasets into two categories for hyperparameter experiments [^4] https://anonymous.4open.science/r/EWGSL-C021/ReadMe.txt/. We search the hyperparameters $\alpha$, $K$, and $\eta$ within the ranges {1,…,2}, {2,…,12}, and {0,…,0.5}, respectively. The results show that the model performs best when $\alpha$, $K$, and $\eta$ are set to {1.5, 1.5}, {6, 8}, and {0.05, 0.1}, respectively. We employ the Adam optimizer to minimize the total loss.
To improve the generalization ability and reduce the impact of randomness on the experimental results, we conducted multiple experiments with different random seeds and averaged the results.
Metrics: We evaluate the accuracy of a node classification by two common metrics Accuracy (ACC) and Micro-F1 (F1) [34], and a higher score means a more accurate result.
@@ IV-B Experment Results
Moreover, although the performance of SGAT is close to that of EWGSL, it does not surpass it. This is primarily because SGAT assigns a unique attention coefficient to each edge, which restricts its ability to capture global information and reduces flexibility in adapting to contextual changes. As a result, SGAT loses some details when processing intricate information.
@@ IV-C Ablation Study
[…]
Moreover, the impact of the graph sparsity structure learning module on the ML-100k__\__ES dataset is smaller compared to other datasets, indicating that its advantages are more pronounced when handling denser graph data. Specifically, the graph structure of the ML-100k__\__ES dataset is relatively sparse, so the advantage of optimizing the graph structure and capturing node relationships is less significant than in denser graph data. In contrast, the node relationships in dense graph data are more complex, and the graph sparsity structure learning approach can improve the graph structure more effectively, thereby enhancing the model’s performance.
@@ IV-D Case Study
The heatmap in Fig. 3 illustrates a comparison of the attention weights calculated by GAT and EWGSL. By comparing the results from both methods, it is clear that EWGSL can more effectively adjust the distribution of attention, enhancing the expressiveness of the graph structure. As shown in Fig. 3 (b), Node 2 has strong connections with two neighboring nodes, and after applying EWGSL, the weights of these two neighbors significantly increase, with their colors deepening, indicating that their importance within the graph is more strongly highlighted (see Fig. 3 (a)). Moreover, EWGSL is also capable of effectively identifying and suppressing the influence of noise. In this experiment, EWGSL successfully identified three irrelevant nodes with weak relationships to Node 2 and its neighbors, effectively reducing their weights and mitigating their interference with the graph representation. In contrast, the traditional GAT performed poorly in handling these irrelevant nodes, failing to effectively distinguish between important nodes and noise. This result demonstrates that EWGSL can better optimize the graph structure when dealing with noisy graph data, thereby improving the model’s performance.

### v2

@@ I Introduction
To enhance the ability of the model to handle noisy edges, we combine two approaches to address node classification in weighted graphs, facing two primary challenges. First, edge weights on weighted graphs can describe fine-grained relationships between nodes, with higher weights indicating stronger connections and greater influence on information propagation[20][21]. However, existing weight learning methods primarily focus on the structural characteristics of the graph and do not adequately account for edge weights. How to learn more fine-grained weights between nodes is a challenge. Second, noisy edges in weighted graphs are characterized by anomalous edge weights. Specific scenarios, such as network failures that affect communication between users for a period of time, can cause edge weights that describe the communication frequency to deviate significantly from normal conditions[22]. How to adaptively correct anomalous edge weights and remove noisy edges is a challenge.
To address these challenges, we propose an Edge Weight-aware Graph Structure Learning (EWGSL) method, a novel model that employs edge weight-aware methods to learn fine-grained weights and employing Graph Structure Learning (GSL)[22] to reduce noise interference. Leveraging the excellent scalability of Graph Attention Networks (GATs)[23], we use GAT to learn node representations. We redefine the calculation of attention coefficients in GAT by incorporating both node features and edge weights to capture the relationships between nodes accurately. We apply GSL to sparsify the attention coefficients, setting the attention weights of irrelevant nodes to zero and concentrating them on relevant nodes to effectively identify noisy edges. Additionally, we modify the noise contrastive estimation (InfoNCE) loss function[24] to dynamically utilize the weights from the denoised graph structure, thereby enhancing node classification performance.
@@ II-A Weight Learning Methods
WFSM-MaxPWS[20] prunes the subgraph expansion process by calculating the “Maximum Possible Weighted Support”, retaining only subgraphs with higher weighted support. This reduces computational overhead and improves mining efficiency while ensuring the completeness of the results. MetaGC[8] adaptively adjusts the weight of each edge in the graph using a meta-model, specifically lowering the weight of noisy edges to mitigate their impact on the clustering results. Additionally, it employs a decomposable clustering loss function and end-to-end training, allowing the method to enhance robustness against noisy edges without relying on external denoising steps.
@@ II-B Structure Learning Methods
Structure learning methods handle noisy edges by optimizing the graph structure, primarily through identifying and removing noisy edges, and reconstructing the graph connectivity. This enhances the propagation of information between similar nodes. This approach ensures a more accurate graph structure, improves the quality of node representations, and reduces the interference of noise on model learning.
@@ III-A Problem Definition
Problem Formulation: Given a weighted graph $\mathrm{G}$, and a small set of nodes $\mathrm{V_{L}}$ with labels $\mathrm{Y_{L}}$, the goal of the node classification task is to predict unlabeled nodes $\mathrm{V_{U}}$ with $\mathrm{Y_{U}}$, where ${\mathrm{V_{L}}}\cup{\mathrm{V_{U}}}=\mathrm{V}$.
@@ III-B Edge-aware Weight Learning
$\vspace{-0.2cm}e_{ij}=\rho_{ij}\cdot{\boldsymbol{a}(\textbf{W}{\vec{h}_{i}},\textbf{W}{\vec{h}_{j}})},\vspace{-0.2cm}$ (1)
@@
$\rho_{ij}=\frac{w_{ij}}{\sum\limits_{k\in{{\mathcal{N}_{i}}}}w_{ik}}\;(i\notin{\mathcal{N}_{i}}),\vspace{-0.2cm}$ (2)
@@ III-C Graph Sparsity Structure Learning
${e}_{ij}^{\prime}=\mathrm{{\alpha}\text{-}entmax}({e_{ij}})={[(\alpha-1){e_{ij}}-\tau]_{+}^{1/{\alpha-1}}},\vspace{-0.2cm}$ (4)
@@
${\vec{h}_{i}}^{\prime}=\mathrm{F}(\Delta{\vec{h}_{i}}^{\prime})=\mathrm{F}(\beta_{k}{\vec{h}_{ki}}^{\prime})\quad(k=1,2,\ldots,K),\vspace{-0.2cm}$ (7)
@@
$\mathcal{H}=[{\vec{h}_{1}}^{\prime},{\vec{h}_{2}}^{\prime},\cdots,{\vec{h}_{n}}^{\prime}],\vspace{-0.2cm}$ (8)
@@ III-D Loss Function
$\mathcal{M}=[\mathrm{softmax}({\vec{h}_{1}}^{\prime}),\mathrm{softmax}({\vec{h}_{2}}^{\prime}),\cdots,\mathrm{softmax}({\vec{h}_{n}})].\vspace{-0.2cm}$ (9)
@@
$\hat{y}_{i}={\mathrm{{\underset{{\textit{k}}\in\{1,2,\cdots,c\}}{argmax}}}m_{ik}}.\vspace{-0.2cm}$ (10)
@@ 13      Backpropagate and update model parameters using $L$;
$\vspace{-0.25cm}L_{I}=-\frac{1}{n}\sum_{i=1}^{n}\log\frac{\lambda_{p}\exp\left({\text{sim}({\vec{h}_{i}}^{\prime},{\vec{h}_{j}}^{\prime})}/{t}\right)}{{\sum\limits_{k\in N_{i}}}\lambda_{n}\exp\left({\text{sim}({\vec{h}_{i}}^{\prime},{\vec{h}_{k}}^{\prime})}/{t}\right)},$ (12)
@@ IV-A Experiment Settings
ML-100k__\__ES [^1] https://github.com/wtt2022/DataSet/ is constructed from the open dataset MovieLens 100K [^2] https://grouplens.org/datasets/movielens/. The dataset includes multiple files, and we focus on u.data and u.item. We will release this dataset.
@@
We use a three-layer GAT to learn node representations, with hidden layers of sizes 256 and 128, and input and output sizes of $n$ and $c$, respectively. The model trains for 100 epochs with a learning rate of 0.005. Based on the dataset construction, we divide the datasets into two categories for hyperparameter experiments. We search the hyperparameters $\alpha$, $K$, and $\eta$ within the ranges {1,…,2}, {2,…,12}, and {0,…,0.5}, respectively. The results show that the model performs best when $\alpha$, $K$, and $\eta$ are set to {1.5, 1.5}, {6, 8}, and {0.05, 0.1}, respectively. We employ the Adam optimizer to minimize the total loss.
To enhance generalization and reduce randomness effects, we conduct multiple experiments with different random seeds and average the results.
Metrics: We evaluate the accuracy of a node classification by two common metrics Accuracy (ACC) and Micro-F1 (F1), and a higher score means a more accurate result.
@@ IV-B Experment Results
Moreover, although the performance of SGAT is close to that of EWGSL, it does not surpass it. This is primarily because SGAT assigns a unique attention coefficient to each edge, limiting its global information capture and adaptability to context changes. As a result, SGAT loses some details when processing intricate information.
@@ IV-C Ablation Study
Moreover, the impact of the graph sparsity structure learning module on the ML-100k__\__ES dataset is smaller compared to other datasets, indicating that its advantages are more pronounced when handling denser graph data. Specifically, the graph structure of the ML-100k__\__ES dataset is relatively sparse, sothe advantage of optimizing graph structure and capturing node relationships is less significant than in denser data. In contrast, the node relationships in dense graph data are more complex, and the graph sparsity structure learning approach can improve the graph structure more effectively, thereby enhancing the performance of the model.
@@ IV-D Case Study
The heatmap in Fig. 3 illustrates a comparison of the attention weights calculated by GAT and EWGSL. By comparing the results from both methods, it is clear that EWGSL can more effectively adjust the distribution of attention, enhancing the expressiveness of the graph structure. As shown in Fig. 3 (b), Node 2 has strong connections with two neighboring nodes, and after applying EWGSL, the weights of these two neighbors significantly increase, with their colors deepening, indicating that their importance within the graph is more strongly highlighted (see Fig. 3 (a)). Moreover, EWGSL is also capable of effectively identifying and suppressing the influence of noise. In this experiment, EWGSL successfully identified three irrelevant nodes with weak relationships to Node 2 and its neighbors, effectively reducing their weights and mitigating their interference with the graph representation. In contrast, the traditional GAT performed poorly in handling these irrelevant nodes, failing to effectively distinguish between important nodes and noise. This result demonstrates that EWGSL can better optimize the graph structure when dealing with noisy graph data, thereby improving the performance of the model.
@@ V Conclusion
After completing the current research, our future research will focus on addressing the challenges posed by the rapid growth of data volume. As the data volume surges sharply, the computational complexity of large weighted graphs increases dramatically, and the traditional computing resources and algorithmic efficiency have revealed significant limitations. Therefore, it is of utmost urgency to optimize existing algorithms and explore new computing architectures to reduce the computational complexity.
Acknowledgment
This research is supported in part by the National Science Foundation of China (No. 62302469), the Natural Science Foundation of Shandong Province (ZR2023QF100, ZR2022QF050).